# Superadiabatic Landau-Zener transitions

### v1

@@ Abstract
Conclusions. We have proposed a generalization of the LZ model that gives rise to a superadiabatic (SA) regime while maintaining the same instantaneous eigenenergies as in the LZ case. This model is relevant, as it can be used to describe, e.g., the valley transitions during electron spin shuttling in SiGe quantum dots. The presented approach reveales how the path of the instantaneous eigenvectors in $xz$ plane determine the transition probability $P$. We have described the SA regime, where $P$ is significantly lower than in the LZ case, and where $P=0$ in the unconditionally adiabatic case, no matter the driving velocity and energy gap. This offers a new path towards engineering driven dynamics with reduced excitation probability in the SA regime. We have also found that LZ transitions are possible even without the presence of any avoided crossing point in the energy levels, revealing that transitions are in fact a consequence of the rotation of the eigenvectors in the Bloch sphere. Future extensions may include dissipative systems which have been widely investigated in the context of LZ transitions [50, 51, 52, 53, 54]. The effects of unavoidable dissipation in the extended model and SA regime deserve further investigation.

### v2

@@ IV Consequences for shuttling
V Conclusions
We have proposed a generalization of the LZ model that gives rise to a superadiabatic (SA) regime while maintaining the same instantaneous eigenenergies as in the LZ case. This model is relevant, as it can be used to describe, e.g., the valley transitions during electron spin shuttling in SiGe quantum dots. The presented approach reveals how the Hamiltonian curve in the $xz$ plane determines the transition probability $P$. We have described the SA regime, where $P$ is significantly lower than in the LZ case, and the unconditionally adiabatic case, where $P=0$ no matter the driving velocity and energy gap. This offers a new path towards the engineering of driven dynamics with reduced excitation probability in the SA regime. We have also found that LZ transitions are possible even without the presence of any avoided crossing point in the energy levels, revealing that transitions are a consequence of the rotation of the eigenvectors in the Bloch sphere. Future extensions may include dissipative systems which have been widely investigated in the context of LZ transitions [51, 52, 53, 54, 55]. The effects of unavoidable dissipation in the extended model and SA regime deserve further investigation. Quantum control schemes for quantum systems coupled to a thermal bath can also be applied to our model [56, 57], particularly to the SA regime. Future work may also explore the influence of the topology of the closed paths of the instantaneous eigenvectors in the adiabatic regime.
@@ Acknowledgements.
Appendix A Valley splitting in a silicon quantum dot
Within the effective mass theory, the two low-lying valley states of a silicon quantum dot in a SiGe/Si/SiGe heterostructure grown in the $\hat{z}$ direction can be written as [58]
|±$|\pm z\rangle=\Psi_{xyz}(\textbf{r})e^{\pm ik_{0}z}u_{\pm z}(\textbf{r}),$ (6)
where $\Psi_{xyz}(\textbf{r})$ is the envelope function, $u_{\pm z}(\textbf{r})$ are the periodic parts of the Bloch wavefunctions at the two conduction band minima (valleys) of silicon, $k_{0}=0.82(2\pi/a_{0})$ is the Bloch wavenumber at the two valleys, and $a_{0}=0.543$ nm is the length of the Si cubic unit cell.
We model the confinement potential due to the SiGe barriers by a sum of delta functions at the location of each Ge atom [15, 16],
$U(x,y,z)=\lambda\sum_{i}\delta(x-x_{i})\delta(y-y_{i})\delta(z-z_{i}),$ (7)
[…]
The intervalley coupling is given by
$\Delta(\textbf{r})$ $=$ $\langle+z|U(x,y,z)|-z\rangle$ (8)
$=$ $C_{0}\int e^{-2ik_{0}z}U(x,y,z)|\Psi_{x,y,z}(\textbf{r})|^{2}d^{3}x,$
where $C_{0}=-0.2607$ results from the periodic parts of the Bloch wavefunctions [59, 60]. The total valley splitting is then given by
$E_{\rm VS}(\textbf{r})=2|\Delta(\textbf{r})|.$ (9)
Due to the random distribution of the Ge atoms at the SiGe/Si interfaces, the so-called alloy disorder, the intervalley coupling, and consequently the valley splitting, depends on the location of the quantum dot, which gives rise to the random oscillations of the valley splitting in Fig. 1 (a) and the random path in the complex plane in Fig. 1 (b) during charge/spin shuttling.
[…]
Appendix B Valley shuttling fidelity
The Hamiltonian for the two valley states can be written as
$H=\left(\begin{array}[]{cc}0&\Delta(\textbf{r})\\ \Delta(\textbf{r})^{*}&0\end{array}\right).$ (12)
Using the polar decomposition $\Delta=|\Delta|e^{i\theta}$, the two valley eigenstates can be written as
$\alpha_{\pm}(\textbf{r})=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}e^{i\theta(\textbf{r})/2}\\ \pm e^{-i\theta(\textbf{r})/2}\end{array}\right).$ (15)
These are the instantaneous eigenstates of Eq. (12) with corresponding instantaneous eigenenergies,
$E_{\pm}(\textbf{r})=\pm|\Delta(\textbf{r})|,$ (16)
giving rise to the valley splitting $E_{\rm VS}(\textbf{r})=E_{+}(\textbf{r})-E_{-}(\textbf{r})$ in Eq. (4). For a quantum dot moving in the $\hat{\textbf{x}}$ direction at the (shuttling) velocity $v$, the Schrödinger equation,
$i\hbar\partial_{t}\psi(t)=H(t)\psi(t),$ (17)
can be rewritten employing a change of variables, $\textbf{r}=v\hat{\textbf{x}}t$, and thus $x=vt$, as
$i\hbar v\partial_{x}\psi(x)=H(x)\psi(x).$ (18)
We solve Eq. (18) numerically considering that initially the electron is in the valley ground state $\alpha_{-}$ and write
|$\ket{\psi(x)}=a(x)\ket{\alpha_{-}(x)}+b(x)\ket{\alpha_{+}(x)}.$ (19)
The probability of a valley transition is then given by
$P(x)=|\bra{\alpha_{+}}(x)\ket{\psi(x)}|^{2}=|b(x)|^{2}.$ (20)